OpenStudy (diana.xl):

A new online store charges all customers one price to download any song and another price to download any game. Jenna paid $20.50 to download eight songs and three games, and Dwayne paid $22.50 to download four songs and five games. How much does the store charge to download each song and game?

OpenStudy (diana.xl):

a. Write a system of linear equations based on the description. Use x to represent the cost of each song and y to represent the cost of each game. b. Solve the system using the elimination method. Show all the steps leading to your answer.

OpenStudy (diana.xl):

@imqwerty

imqwerty (imqwerty):

let the price to download 1 song be x let the price to download 1 game by y so 8x+3y-20.5 and 4x+5y=22.5 can u solve get the answers now?

OpenStudy (diana.xl):

@MrNood

imqwerty (imqwerty):

so now we have the equations and we have to solve them using elimination method

imqwerty (imqwerty):

first -> I made a typo its not 8x+3y-20.5 ...its this--> 8x+3y=20.5

imqwerty (imqwerty):

8x+3y=20.5 -eq 1 and 4x+5y=22.5 -eq 2 we multiply the 2nd equation by 2 we get this- 2(4x+5y)=2(22.5) 8x+10y=45 -eq 3 now we subtract the eq 3 from eq 1 we get- 8x+3y-8x-10y=20.5-45 -7y=-24.5 y=3.5 we know 4x+5y=22.5 put y=3.5 to get x 4x+5(3.5)=22.5 x=1.25

OpenStudy (diana.xl):

http://prntscr.com/8t8zx7 like this?